OpenStudy (anonymous):

i've been stuck on this for days and will fail if i don't complete this assignment. i have tried it twice and failed.

OpenStudy (aum):

First list the initial stock price for each of the stocks: A, B and C. Next, find the rate of growth of each of the stocks: A, B and C.

OpenStudy (aum):

Initial stock prize is when x = 0.

OpenStudy (aum):

List initial stock prices for each of the stock: Stock A: \(f(x) = 25(1.08)^x\). When x = 0, f(x) = 25 Stock B: When x = 0, f(x) = 22 Stock C: The initial price of this stock is $30. Stock D should have the second lowest starting price. Can you pick a number that is the second lowest based on the above data?

OpenStudy (anonymous):

24?

OpenStudy (aum):

That will work. Stock D initial stock price: $24. Next, determine the rates of growth for each stock because for stock D has to have the second highest rate of growth.

OpenStudy (aum):

Stock A: \(f(x) = 25(1.08)^x = 25(1+0.08)^x\). The rate of growth for this stock is 0.08 or 8%. Stock B: Decreases in value. So we can treat this as the lowest rate of growth because it is negative rate of growth. We don't need the exact number. Stock C: Increases by 4%. Let us put them in ascending order: negative growth, 4%, 8%. Pick a rate of growth for stock D that is the second HIGHEST rate of growth.

OpenStudy (aum):

Okay, that will work. Stock D: Initial price $24. Rate of growth 6%. A suitable function for stock D is: \(f(x) = 24(1+0.06)^x = 24(1.06)^x\). Stock D: \(f(x) = 24(1.06)^x\).

OpenStudy (aum):

Yes. You can provide the function and explain in sentences that the initial price of this stock is $24 which is the second lowest amongst stocks A, B, C and D. The rate of growth for stock D is 6% which is the second highest amongst stocks A, B, C and D.

OpenStudy (aum):

Graph the price function for Stock D: \(f(x) = 24(1.06)^x\)  Key features: y-axis is the price; x axis is number of days. The function is \(f(x) = 24(1.06)^x\). It crosses the y-axis at (0,24) which implies that when x = 0, y = 24 or the initial stock price is $24. The function is an exponential growth function and the rate of growth is 6%. The curve to the left of the y-axis can be ignored as the domain of the function is \(x \ge 0\) and the range of the function is \(y \ge 24\).

OpenStudy (aum):

#3. Stock D’s new price function, f(x) + g(x) = \(24(1.06)^x - 6\). This new price function will shift the graph of the f(x) function DOWN by 6 units.